OpenStudy (anonymous):

The formula for degrees Celsius is: C = five over nine (F − 32), where F stands for degrees Fahrenheit. Which of the following equations shows the formula correctly solved for F? F = five over nine C + 32 F = nine over five C + 32 F = five over nine C − 32 F = nine over five C − 32

OpenStudy (lalaly):

the second

OpenStudy (anonymous):

C = 5/9* (F -32) 9/5*C = F -32 9/5 * C + 32 =F Second option